OpenStudy (itsonlycdeee):

Find the derivative of: f(x)=[ x^(1/2)+1]/[x^2 + 1]

zepdrix (zepdrix):

\[\Large\bf\sf f(x)\quad=\quad \frac{x^{1/2}+1}{x^2+1}\] We'll need to apply the Quotient Rule:\[\Large\bf\sf f'(x)\quad=\quad \frac{\color{royalblue}{(x^{1/2}+1)'}(x^2+1)-(x^{1/2}+1)\color{royalblue}{(x^2+1)'}}{(x^2+1)^2}\]

zepdrix (zepdrix):

Does the setup make sense? We need the derivatives of the blue parts.

OpenStudy (itsonlycdeee):

Yes, the setup makes sense. So it would become [1/2x^{-1/2}(x^2+1) - x^(1/2)+1)(2x)] / (x^2+1)^2 right?

zepdrix (zepdrix):

Mmm ya looks good!

OpenStudy (itsonlycdeee):

Ok, thank you. But I'm on confused on what to do after that lol.

OpenStudy (itsonlycdeee):

The answer in the book is: (0.5x^{-1/2})[x^2+1-4x^{3/2}(x^{1/2})]/(x^2+1)^2 OR -3x^2-4x^{3/2} + 1 / 2x^{1/2}(x^2+1)^2

zepdrix (zepdrix):

So the first answer, it looks like all they did was: factor a (1/2)x^(-1/2) out of each term.

zepdrix (zepdrix):

\[\Large\bf\sf f'(x)\quad=\quad \frac{\color{orangered}{(\frac{1}{2}x^{-1/2})}(x^2+1)-(x^{1/2}+1)\color{orangered}{(2x)}}{(x^2+1)^2}\]

zepdrix (zepdrix):

So if we factor that first orange part out of each term.. we should get uhhhh.. So it will simply disappear from the first term, yes? Then we need to alter the 2x to compensate.

zepdrix (zepdrix):

It doesn't really matter though, dont' trust your book answers too much... They tend to oversimplify things D:

zepdrix (zepdrix):

\[\Large\bf\sf f'(x)\quad=\quad \color{orangered}{\left(\frac{1}{2}x^{-1/2}\right)}\frac{(x^2+1)-(x^{1/2}+1)\left(\frac{2x}{\left(\dfrac{1}{2}x^{-1/2}\right)}\right)}{(x^2+1)^2}\]

OpenStudy (itsonlycdeee):

Okay then aha. Thank you!

zepdrix (zepdrix):

Mmm I made that look a bit ugly :P So dividing 2x and (1/2)x^(-1/2) gives us:\[\Large\bf\sf f'(x)\quad=\quad \color{orangered}{\left(\frac{1}{2}x^{-1/2}\right)}\frac{(x^2+1)-(x^{1/2}+1)\left(4x^{3/2}\right)}{(x^2+1)^2}\]

zepdrix (zepdrix):

And then that's a little bit closer to what they have right? :\ blah lol

OpenStudy (itsonlycdeee):

I think lol. The simplified ones the book has gets confusing.

zepdrix (zepdrix):

ya :c